6. DDT, an insecticide harmful to fish, birds and humans is produced by the following reaction

2C6H5CL + C2HOCL3 ———–> C14H9Cl5 + H2O

Chlorobenzene Chloral DDT

In a government lab, 1142 g of chlorobenzene is reacted with 485 g of chloral. (a) What mass of DDT is formed, assuming 100% yield

(b) Which reactant is limiting and in excess? Justify your answer.

(c) What mass of excess reactant is left over

(d) If the actual yield of DDT is 200 g, what is the percentage yield


1
Expert's answer
2021-03-10T06:52:36-0500



2C6H5Cl + C2HOCl3 ———–> C14H9Cl5 + H2O

The limiting reactant can be determined via calculating the number of moles:

n(chlorobenzene) = m(chlorobenzene) / [2 × Mr(chlorobenzene)] = 1142 g / [2 × 112.56 g/mol] = 5.07 mol

n(chloral) = m(chloral) / Mr(chloral) = 485 g / 147.388 g/mol = 3.29 mol

a) Since chloral is a limiting reactant, the mass of DTT is:

m(DTT) = n(cloral) × Mr(DTT) = 3.29 mol × 154.253 g/mol = 507.5 g

b) Chloral is a limiting reactant.

c) Mass of chlorobenzene used in the reaction:

m(chlorobenzene) = n(chloral) × 2 × Mr(chlorobenzene) = 3.29 mol × 2 × 112.56 g/mol = 740.65 g

Mass of chlorobenzene left:

m(chlorobenzene) = 1142 g - 740.65 g = 401.35 g

d) The percentage yield is:

%w = (actual yield / theoretical yield) × 100% = (200 g / 507.5 g) × 100% = 39.4%